OpenStudy (anonymous):

Input the equation of the given line in standard form. The line with m = -3/4 and b = -2.

OpenStudy (anonymous):

@Rainbow_Dash @Hero

OpenStudy (rainbow_dash):

standard form. is this slope intercept form or point slope form?

OpenStudy (rainbow_dash):

standard is going to be y=-3/4x-2, point slope is going to be y-(y_1)=-3/4(x-x_1), y_1 and x_1 being a single point on the line.

OpenStudy (anonymous):

no idea! thats my problem

OpenStudy (aroub):

I looked it up and it says => AX+BY=C is the standard form

OpenStudy (rainbow_dash):

ahh. my bad.

OpenStudy (rainbow_dash):

and bys tandard in my 2nd post i meant y-intercept* not standard

OpenStudy (anonymous):

thanx guys!

OpenStudy (rainbow_dash):

not a problem.